# Why Agentic Theorem Prover Works: A Statistical Provability Theory of Mathematical Reasoning Models

### v1

@@ 1 Introduction
Large language models (LLMs) have rapidly improved at mathematical reasoning (Barreto, 2026; Bryan et al., 2026; Epoch AI, 2026), and a parallel effort has produced increasingly capable agentic theorem provers: systems that combine an LLM policy, a planner for decomposition, a retriever over large libraries (such as Mathlib), and a proof assistant (such as Lean, Rocq, and Isabelle, often with SAT/SMT support) that provides rigorous feedback. Despite striking empirical progress, these pipelines are complex and highly modular, and it remains difficult to explain which components affect how—and in what sense they can overcome, or fail to overcome, the classical hardness barriers surrounding proof search. This gap between practice and theory is now a bottleneck: many design choices are guided by heuristics and ablations, while a principled understanding of provability under realistic distributions and finite compute is still missing.
@@ 1.1 Our approach: provability as a finite-horizon reachability problem
To make such statements precise, we model an agentic theorem prover as a finite-horizon Markov decision process (MDP). The state encodes the current proof obligation and verifier feedback (e.g., goals, subgoals, error messages), the action encodes the next proof step (tactic/lemma invocation, parameters, retrieval prompt, decoding controls), and the transition is induced by the proof assistant’s execution and checking. The evaluation target is time-bounded success: reach the solved state within a budget of $B$ interaction steps (or verifier calls). This leads to a canonical object:
  • the provability of an instance $x_{0}$ under budget $B$,
    $V_{B}^{*}(x_{0}):=\sup_{\pi}\Pr^{\pi}(\exists t\leq B:X_{t}\in G\mid X_{0}=x_{0}),$
@@ 1.2 Why Bellman structure matters: certificates and component-wise effects
1.3 From heuristics to qualitative theory: what each pipeline component buys you
This framework lets us reason qualitatively (and, under assumptions, quantitatively) about the impact of common pipeline components:
  • Verifier feedback shrinks the reachable region, eliminates invalid branches early, and can increase action separations; in the theory this improves effective dimension, margins, and certificate tightness.
  • Retrieval over libraries (RAG) reduces approximation error by exposing relevant lemmas/proof schemas; in the theory this is a decrease in $\epsilon_{\mathrm{app}}$ and often a reduction of the effective horizon by enabling shortcuts.
  • Test-time search (beam/rollout) trades compute for probability mass on good trajectories; its benefit is controlled by margin tails and by how quickly near-optimal actions remain in top-$k$ sets.
  • Representation geometry (how proof states are embedded into a metric space) controls statistical rates through covering numbers / metric entropy; in favorable geometries, estimation error decays much faster with the number of verifier interactions.
@@ 2 Problem Formulation
Here we formalize interactive formal theorem proving (with an LLM agent and an exact verifier/proof assistant) as a time-bounded reachability problem in a Markov decision process (MDP). We then develop a Bellman-structure–based evaluation of (i) provability (optimal success probability), (ii) deviation from optimality of score-based planning (greedy/top-kk/beam/rollout), and (iii) certificate-style upper/lower bounds. Finally, we discuss easy vs hard instances.
@@
$(K,d_{K})$ denotes a compact metric space used as an embedding space for individual goals, $M(K)$ denote the set of finite positive Borel measures on $K$. For $W>0$:
$M_{\leq W}(K):=\{\mu\in M(K):\mu(K)\leq W\}.$
For a bounded function $f:K\to\mathbb{R}$:
$\|f\|_{\infty}:=\sup_{x\in K}|f(x)|,\quad\mathrm{Lip}(f):=\sup_{x\neq y}\frac{|f(x)-f(y)|}{d_{K}(x,y)},$
$\|f\|_{\mathrm{BL}}:=\max\{\|f\|_{\infty},\mathrm{Lip}(f)\}.$
The bounded Lipschitz metric on $M(K)$ is
$d_{\mathrm{BL}}(\mu,\nu):=\sup_{\|f\|_{\mathrm{BL}}\leq 1}\int_{K}fd(\mu-\nu).$
“Goal solved” is represented by the zero measure $0\in M(K)$.
@@ 3.2 Measure-valued encoding of “a finite set of goals”
Let $\mathcal{G}_{0}$ denote the (discrete) set of syntactic goals (or goal+context objects). Fix an embedding map
$\phi:\mathcal{G}_{0}\to K.$
If a proof state $x$ contains goals $g_{1},\dots,g_{n}\in\mathcal{G}_{0}$ (allowing multiplicity), define the goal measure
$\mu(x):=\sum_{i=1}^{n}\delta_{\phi(g_{i})}\in M(K).$
We note: (1) The solved state (“no goals”) corresponds to $\mu(x)=0$. (2) The number of goals is the total mass: $\mu(x)(K)=n$. (3) If we enforce (or truncate to) $\mu(x)(K)\leq W$, then $\mu(x)\in M_{\leq W}(K)$.
@@
If $(K,d_{K})$ is compact, then $(M_{\leq W}(K),d_{\mathrm{BL}})$ is a compact metric space.
@@ 3.3 MDP
State space: $\mathcal{X}:=M_{\leq W}(K)$ (or a truncated version; see Section˜8).
@@
Transition kernel: $P(\cdot\mid\mu,a)$ is a Borel probability measure on $\mathcal{X}$.
@@ 3.4 Provability as a time-bounded reachability objective
Fix a horizon $B\in\mathbb{N}$. For a policy $\pi$ and initial state $\mu\in\mathcal{X}$, define the time-bounded success probability
$V_{B}^{\pi}(\mu):=\Pr^{\pi}_{\mu}\left(\exists t\in\{0,1,\dots,B\}\ \text{such that}\ \mu_{t}\in G\right).$
@@
$V_{b}^{*}(\mu):=\sup_{\pi}\Pr^{\pi}_{\mu}\left(\exists t\in\{0,1,\dots,b\}\ \text{s.t.}\ \mu_{t}\in G\right).$
Then $V_{0}^{*}(\mu)=\mathbf{1}[\mu\in G]$.
@@
$Q_{b}^{*}(\mu,a):=\int_{\mathcal{X}}V_{b-1}^{*}(\mu^{\prime})P(d\mu^{\prime}\mid\mu,a).$
@@ 4 Existence of optimal value functions and optimal Markov policies
$(\mu,a)\mapsto\int_{\mathcal{X}}f(\mu^{\prime})P(d\mu^{\prime}\mid\mu,a)$
@@
This implies that $(\mu_{n},a_{n})\to(\mu,a)$ entails $P(\cdot\mid\mu_{n},a_{n})\Rightarrow P(\cdot\mid\mu,a)$ weakly.
@@
Assume: (1) $\mathcal{X}=M_{\leq W}(K)$ is compact under $d_{\mathrm{BL}}$ (Theorem˜1), (2) $\mathcal{A}$ is compact, (3) the kernel $P$ satisfies Assumption˜1, (4) $G=\{0\}$ is closed (hence $\mathbf{1}[\mu\in G]$ is upper semicontinuous).
Then for each $b=1,\dots,B$: (a) The Bellman recursion defines a Borel-measurable $V_{b}^{*}$, (b) For each $\mu$, the supremum $\sup_{a\in\mathcal{A}}Q_{b}^{*}(\mu,a)$ is attained by some $a_{b}^{*}(\mu)\in\mathcal{A}$, (c) hence there exists an optimal deterministic Markov policy $\pi^{*}=(a_{b}^{*})_{b=1}^{B}$ achieving $V_{B}^{*}$.
@@ 5 Bellman inequalities and provability certificates
$(TV)(\mu):=\mathbf{1}[\mu\in G]\ \vee\ \sup_{a\in\mathcal{A}}\int_{\mathcal{X}}V(\mu^{\prime})P(d\mu^{\prime}\mid\mu,a).$
@@
For an initial state $\mu_{0}$, any pair $(L_{B}(\mu_{0}),U_{B}(\mu_{0}))$ such that
$L_{B}(\mu_{0})\leq V_{B}^{*}(\mu_{0})\leq U_{B}(\mu_{0})$
@@ 6.1 Score functions and induced policies
$\pi_{h}(\mu,b)\in\arg\max_{a\in\mathcal{A}}h_{b}(\mu,a).$
Similarly, define top-$k$ candidate sets $\mathrm{Top}\text{-}k(h_{b}(\mu,\cdot))$ and beam-search variants (beam introduces a tree policy, but for analysis we track how often near-optimal actions remain in the candidate set).
@@ 6.2 Uniform score approximation and worst-case regret
Under Assumption˜2, for every $\mu$ such that $(\mu,a)\in\mathcal{D}$ for all relevant $a$,
$\sup_{a}Q_{b}^{*}(\mu,a)-Q_{b}^{*}(\mu,\pi_{h}(\mu,b))\leq 2\varepsilon_{b}.$
@@
Let $a^{*}\in\arg\max_{a}Q_{b}^{*}(\mu,a)$ and $\hat{a}\in\arg\max_{a}h_{b}(\mu,a)$. Then $Q_{b}^{*}(\mu,a^{*})\leq h_{b}(\mu,a^{*})+\varepsilon_{b}\leq h_{b}(\mu,\hat{a})+\varepsilon_{b}\leq Q_{b}^{*}(\mu,\hat{a})+2\varepsilon_{b}$. Rearrange. ∎
@@
Under Assumption˜2, for any initial $\mu_{0}$ whose relevant trajectories remain inside $\mathcal{D}$,
$0\leq V_{B}^{*}(\mu_{0})-V_{B}^{\pi_{h}}(\mu_{0})\leq 2\sum_{b=1}^{B}\varepsilon_{b}.$
@@ 6.3 Certificate construction from score bounds
$U_{b}(\mu):=\mathbf{1}[\mu\in G]\ \vee\ \mathrm{clip}\Big(\sup_{a}h_{b}(\mu,a)+\varepsilon_{b}\Big),$
$L_{b}(\mu):=\mathbf{1}[\mu\in G]\ \vee\ \mathrm{clip}\Big(\sup_{a}h_{b}(\mu,a)-\varepsilon_{b}\Big),$
@@
For each $\mu$, $\sup_{a}Q_{b}^{*}(\mu,a)\in[\sup_{a}h_{b}(\mu,a)-\varepsilon_{b},\ \sup_{a}h_{b}(\mu,a)+\varepsilon_{b}]$, then apply Theorem˜3. ∎
@@ 6.4 Top-kk/beam and a margin condition (fast rates)
$\Delta_{b}^{(k)}(\mu):=Q_{b}^{*}(\mu,a_{1}^{*}(\mu))-Q_{b}^{*}(\mu,a_{k+1}^{*}(\mu)),$
where $a_{i}^{*}(\mu)$ are actions sorted by $Q_{b}^{*}(\mu,\cdot)$.
@@
$\nu_{b}(\Delta_{b}^{(k)}(\mu)\leq t)\leq C_{\Delta}t^{\beta}.$
@@
$\mathbb{E}_{\mu_{0}\sim Q}\big[V_{B}^{*}(\mu_{0})-V_{B}^{\pi_{h,k}}(\mu_{0})\big]\ \leq\ C\sum_{b=1}^{B}\varepsilon_{b}^{\beta+1}$
@@
If $\Delta_{b}^{(k)}(\mu)>2\varepsilon_{b}$, uniform perturbation of size $\varepsilon_{b}$ cannot push the best action below rank $k$, so top-$k$ contains an optimal action. Errors only occur on $\Delta_{b}^{(k)}\leq 2\varepsilon_{b}$}, which has probability $\leq C_{\Delta}(2\varepsilon_{b})^{\beta}$, and the loss there is $O(\varepsilon_{b})$. Multiply and sum over $b$. ∎
@@ 7.1 Assumptions
Assume each hypothesis class $\mathcal{H}_{b}$ is uniformly Lipschitz on $\mathcal{D}$: $\mathcal{H}_{b}\subseteq\mathrm{Lip}_{L_{H}}(\mathcal{D},[0,1])$.
@@
$\epsilon_{\mathrm{app},b}:=\inf_{h\in\mathcal{H}_{b}}\ \sup_{(\mu,a)\in\mathcal{D}}|h(\mu,a)-Q_{b}^{*}(\mu,a)|.$
This is a modeling term: it depends on the representation and architecture. Offline library pretraining and richer encoders can be viewed as reducing $\epsilon_{\mathrm{app},b}$ (or localizing the effective hypothesis class).
@@ 7.2 Statistical estimation error (high probability)
$\sup_{f\in\mathcal{F}}\left|\frac{1}{n}\sum_{t=1}^{n}\Big(f(Z_{t})-\mathbb{E}[f(Z_{t})\mid\mathcal{F}_{t-1}]\Big)\right|$
$\quad\leq 2\varepsilon+\sqrt{\frac{2\log\mathcal{N}(\varepsilon;\mathcal{F},\|\cdot\|_{\infty})+2\log(2/\delta)}{n}}.$
@@
Discretize $\mathcal{D}$ by an $\eta$-net and quantize function values; Lipschitzness controls interpolation error. Standard. ∎
@@
Suppose that for each $(\mu,a)\in\mathcal{D}$ we can produce bounded targets $Y\in[0,1]$ such that
$\mathbb{E}[Y\mid\mu,a]=Q_{b}^{*}(\mu,a),$
@@
$\sup_{(\mu,a)\in\mathcal{D}}|\hat{h}_{b}(\mu,a)-Q_{b}^{*}(\mu,a)|\leq\epsilon_{\mathrm{app},b}+\varepsilon_{\mathrm{stat}}(n,\delta),$
@@
$\varepsilon_{\mathrm{stat}}(n,\delta)=O\left(L_{H}^{\frac{d_{\mathcal{D}}}{d_{\mathcal{D}}+2}}\left(\frac{\log(n/\delta)}{n}\right)^{\frac{1}{d_{\mathcal{D}}+2}}\right).$

### v2

@@ 1 Introduction
Large language models (LLMs) have rapidly improved at mathematical reasoning (Barreto, 2026; Bryan et al., 2026; Epoch AI, 2026), and a parallel effort has produced increasingly capable agentic theorem provers: systems that combine an LLM policy, a planner for decomposition, a retriever over large libraries (such as Mathlib), and a proof assistant (such as Lean, Rocq, and Isabelle, often with SAT/SMT solvers) that provides rigorous feedback. Despite striking empirical progress, these pipelines are complex and highly modular, and it remains difficult to explain which components affect how—and in what sense they can overcome, or fail to overcome, the classical hardness barriers surrounding proof search. This gap between practice and theory is now a bottleneck: many design choices are guided by heuristics and ablations, while a principled understanding of provability under realistic distributions and finite compute is still missing.
@@ 1.1 Our approach: provability as a finite-horizon reachability problem
To make such statements precise, we model an agentic theorem prover as a finite-horizon Markov decision process (MDP). The state encodes the current proof obligation and verifier feedback (e.g., goals, subgoals, error messages), the action encodes the next proof step (tactic/lemma invocation, parameters, retrieval prompt, decoding controls), and the transition is induced by the proof assistant’s execution and checking. The evaluation target is time-bounded success: reach the solved state within a budget of $B$ interaction steps (or verifier calls). This leads to a canonical object: the provability of an instance $x_{0}$ under budget $B$,
$V_{B}^{\pi}(x_{0}):=\mathbb{P}^{\pi}(\exists t\leq B:X_{t}\in G\mid X_{0}=x_{0})\quad\text{and}\quad V_{B}^{*}(x_{0}):=\sup_{\pi}V_{B}^{\pi}(x_{0})$
@@ 2 Problem Formulation
We formalize interactive formal theorem proving (with an LLM agent and an exact verifier/proof assistant) as a time-bounded reachability problem in a Markov decision process (MDP). We then develop a Bellman-structure–based evaluation of (i) provability (optimal success probability), (ii) deviation from optimality of score-based planning (greedy/top-kk/beam/rollout), and (iii) certificate-style upper/lower bounds.
@@
Let $(\mathcal{G},d_{\mathcal{G}})$ denote a compact metric space used as an embedding space for individual goals, and let $M(\mathcal{G})$ denote the set of finite positive Borel measures on $\mathcal{G}$. For $W>0$, put
$M_{\leq W}(\mathcal{G}):=\{\mu\in M(\mathcal{G}):\mu(\mathcal{G})\leq W\}.$
For a bounded function $f:\mathcal{G}\to\mathbb{R}$, write
$\|f\|_{\infty}:=\sup_{x\in\mathcal{G}}|f(x)|,\quad\operatorname{Lip}(f):=\sup_{x\neq y}\frac{|f(x)-f(y)|}{d_{\mathcal{G}}(x,y)},\quad\|f\|_{\mathrm{BL}}:=\max\{\|f\|_{\infty},\operatorname{Lip}(f)\}.$
The bounded Lipschitz metric on $M(\mathcal{G})$ is
$d_{\mathrm{BL}}(\mu,\mu^{\prime}):=\sup_{\|f\|_{\mathrm{BL}}\leq 1}\int_{\mathcal{G}}fd(\mu-\mu^{\prime}).$
“Goal solved” is represented by the zero measure $0\in M(\mathcal{G})$.
@@ 3.2 Measure-valued encoding of “a finite set of goals”
Let $\tilde{\mathcal{G}}$ denote the (discrete) set of syntactic goals (or goal+context objects). Fix an embedding map
$\phi:\tilde{\mathcal{G}}\to\mathcal{G}.$
If a proof state $x$ contains goals $\{g_{1},\dots,g_{n}\}\subset\tilde{\mathcal{G}}$ (allowing multiplicity), define the goal measure
$x:=\sum_{i=1}^{n}\delta_{\phi(g_{i})}\in M(\mathcal{G}).$
We note: (1) We use symbol $x$, rather than $\mu$, to mention the goal measure. (2) The solved state (“no goals”) corresponds to $x=0$ (zero-measure). (3) The number of goals is the total mass: $x(\mathcal{G})=n$. (4) If we enforce (or truncate to) $x(\mathcal{G})\leq W$, then $x\in M_{\leq W}(\mathcal{G})$.
@@
If $(\mathcal{G},d_{\mathcal{G}})$ is compact, then $(M_{\leq W}(\mathcal{G}),d_{\mathrm{BL}})$ is a compact metric space.
@@ 3.3 MDP
State space: $\mathcal{X}:=M_{\leq W}(\mathcal{G})$ (or a truncated version; see Section˜8).
@@
Transition kernel: $P(\cdot\mid x,a)$ is a Borel probability measure on $\mathcal{X}$.
@@ 3.4 Provability as a time-bounded reachability objective
Fix a horizon $B\in\mathbb{N}$. For a policy $\pi$ and initial state $x\in\mathcal{X}$, define the time-bounded success probability
$V_{B}^{\pi}(x):=\mathbb{P}^{\pi}\left(\exists t\in\{0,1,\dots,B\}\ \text{such that}\ X_{t}\in G\mid X_{0}=x\right).$
@@
$V_{b}^{*}(x):=\sup_{\pi}\mathbb{P}^{\pi}\left(\exists t\in\{0,1,\dots,b\}\ \text{such that}\ X_{t}\in G\mid X_{0}=x\right).$
Then $V_{0}^{*}(x)=\mathbf{1}[x\in G]$.
@@
$Q_{b}^{*}(x,a):=\int_{\mathcal{X}}V_{b-1}^{*}(x^{\prime})P(dx^{\prime}\mid x,a).$
@@ 4 Existence of optimal value functions and optimal Markov policies
$(x,a)\mapsto\int_{\mathcal{X}}f(x^{\prime})P(dx^{\prime}\mid x,a)$
@@
This implies that $(x_{n},a_{n})\to(x,a)$ entails $P(\cdot\mid x_{n},a_{n})\Rightarrow P(\cdot\mid x,a)$ weakly.
@@
Assume: (1) $\mathcal{X}=M_{\leq W}(\mathcal{G})$ is compact under $d_{\mathrm{BL}}$ (Theorem˜1), (2) $\mathcal{A}$ is compact, (3) the kernel $P$ satisfies Assumption˜1, (4) $G=\{0\}$ is closed (hence $\mathbf{1}[x\in G]$ is upper semicontinuous).
Then for each $b=1,\dots,B$: (a) The Bellman recursion defines a Borel-measurable $V_{b}^{*}$, (b) For each $x$, the supremum $\sup_{a\in\mathcal{A}}Q_{b}^{*}(x,a)$ is attained by some $a_{b}^{*}(x)\in\mathcal{A}$, (c) hence there exists an optimal deterministic Markov policy $\pi^{*}=(a_{b}^{*})_{b=1}^{B}$ achieving $V_{B}^{*}$.
@@ 5 Bellman inequalities and provability certificates
$(TV)(x):=\mathbf{1}[x\in G]\ \vee\ \sup_{a\in\mathcal{A}}\int_{\mathcal{X}}V(x^{\prime})P(dx^{\prime}\mid x,a).$
@@
For an initial state $x_{0}$, any pair $(L_{B}(x_{0}),U_{B}(x_{0}))$ such that
$L_{B}(x_{0})\leq V_{B}^{*}(x_{0})\leq U_{B}(x_{0})$
@@ 6.1 Score functions and induced policies
$\pi_{h}(x,b)\in\arg\max_{a\in\mathcal{A}}h_{b}(x,a).$
Similarly, define top-$k$ candidate sets $\mathrm{Top}\text{-}k(h_{b}(x,\cdot))$ and beam-search variants (beam introduces a tree policy, but for analysis we track how often near-optimal actions remain in the candidate set).
@@ 6.2 Uniform score approximation and worst-case regret
Under Assumption˜2, for every $x$ such that $(x,a)\in\mathcal{D}$ for all relevant $a$,
$\sup_{a}Q_{b}^{*}(x,a)-Q_{b}^{*}(x,\pi_{h}(x,b))\leq 2\varepsilon_{b}.$
@@
Let $a^{*}\in\arg\max_{a}Q_{b}^{*}(x,a)$ and $\hat{a}\in\arg\max_{a}h_{b}(x,a)$. Then $Q_{b}^{*}(x,a^{*})\leq h_{b}(x,a^{*})+\varepsilon_{b}\leq h_{b}(x,\hat{a})+\varepsilon_{b}\leq Q_{b}^{*}(x,\hat{a})+2\varepsilon_{b}$. Rearrange. ∎
@@
Under Assumption˜2, for any initial $x_{0}$ whose relevant trajectories remain inside $\mathcal{D}$,
$0\leq V_{B}^{*}(x_{0})-V_{B}^{\pi_{h}}(x_{0})\leq 2\sum_{b=1}^{B}\varepsilon_{b}.$
@@ 6.3 Certificate construction from score bounds
$U_{b}(x):=\mathbf{1}[x\in G]\ \vee\ \mathrm{clip}\Big(\sup_{a}h_{b}(x,a)+\varepsilon_{b}\Big),$
$L_{b}(x):=\mathbf{1}[x\in G]\ \vee\ \mathrm{clip}\Big(\sup_{a}h_{b}(x,a)-\varepsilon_{b}\Big),$
@@
For each $x$, $\sup_{a}Q_{b}^{*}(x,a)\in[\sup_{a}h_{b}(x,a)-\varepsilon_{b},\ \sup_{a}h_{b}(x,a)+\varepsilon_{b}]$, then apply Theorem˜3. ∎
@@ 6.4 Top-kk/beam and a margin condition (fast rates)
$\Delta_{b}^{(k)}(x):=Q_{b}^{*}(x,a_{1}^{*}(x))-Q_{b}^{*}(x,a_{k+1}^{*}(x)),$
where $a_{i}^{*}(x)$ are actions sorted by $Q_{b}^{*}(x,\cdot)$.
@@
$\nu_{b}(\Delta_{b}^{(k)}(x)\leq t)\leq C_{\Delta}t^{\beta}.$
@@
$\mathbb{E}_{x_{0}\sim Q}\big[V_{B}^{*}(x_{0})-V_{B}^{\pi_{h,k}}(x_{0})\big]\ \leq\ C\sum_{b=1}^{B}\varepsilon_{b}^{\beta+1}$
@@
If $\Delta_{b}^{(k)}(x)>2\varepsilon_{b}$, uniform perturbation of size $\varepsilon_{b}$ cannot push the best action below rank $k$, so top-$k$ contains an optimal action. Errors only occur on $\Delta_{b}^{(k)}\leq 2\varepsilon_{b}$}, which has probability $\leq C_{\Delta}(2\varepsilon_{b})^{\beta}$, and the loss there is $O(\varepsilon_{b})$. Multiply and sum over $b$. ∎
@@ 7.1 Assumptions
Assume each hypothesis class $\mathcal{H}_{b}$ is uniformly Lipschitz on $\mathcal{D}$: $\mathcal{H}_{b}\subseteq\operatorname{Lip}_{L_{H}}(\mathcal{D},[0,1])$.
@@
$\varepsilon_{\mathrm{app},b}:=\inf_{h\in\mathcal{H}_{b}}\ \sup_{(x,a)\in\mathcal{D}}|h(x,a)-Q_{b}^{*}(x,a)|.$
This is a modeling term: it depends on the representation and architecture. Offline library pretraining and richer encoders can be viewed as reducing $\varepsilon_{\mathrm{app},b}$ (or localizing the effective hypothesis class).
@@ 7.2 Statistical estimation error (high probability)
$\sup_{f\in\mathcal{F}}\left|\frac{1}{n}\sum_{t=1}^{n}\Big(f(Z_{t})-\mathbb{E}[f(Z_{t})\mid\mathcal{F}_{t-1}]\Big)\right|\ \leq\ 2\varepsilon+\sqrt{\frac{2\log\mathcal{N}(\varepsilon;\mathcal{F},\|\cdot\|_{\infty})+2\log(2/\delta)}{n}}.$
@@
Discretize $\mathcal{D}$ by an $\eta$-net and quantize function values; Lipschitzness controls interpolation error. ∎
@@
Suppose that for each $(x,a)\in\mathcal{D}$ we can produce bounded targets $Y\in[0,1]$ such that
$\mathbb{E}[Y\mid x,a]=Q_{b}^{*}(x,a),$
@@
$\sup_{(x,a)\in\mathcal{D}}|\hat{h}_{b}(x,a)-Q_{b}^{*}(x,a)|\leq\varepsilon_{\mathrm{app},b}+\varepsilon_{\mathrm{est}}(n,\delta),$
@@
$\varepsilon_{\mathrm{est}}(n,\delta)=O\left(L_{H}^{\frac{d_{\mathcal{D}}}{d_{\mathcal{D}}+2}}\left(\frac{\log(n/\delta)}{n}\right)^{\frac{1}{d_{\mathcal{D}}+2}}\right).$